The price of all houses in Bukit Katil has a mean of RM109,000 and a standard

deviation of RM5,700. Find the probability that the mean price of a random sample of

75 selected houses would be

i. within RM2,200 of the population mean.

ii. more than the population mean by RM1,000 or more.

iii. between RM107,800 and RM110,100.


1
Expert's answer
2021-02-10T01:48:03-0500

The price of all houses in Bukit Katil has a mean RM "109000" and standard deviation RM "5700" .

Let "X" be a random variable defines mean price of "75" selected houses

Then "\mu =109000" and "\sigma =5700" .

Let "Z=\frac{X-\mu}{\sigma}" .Then "Z=\frac{X-109000}{5700}".

(i) We have to find the probability of mean price of "75" selected houses would be within RM "2200" .

"\therefore X=(109000+2200)=111200"

Now "P(X<111200)=P(Z<\frac{111200-109000}{5700})"

"=P(Z<0.39)"

"=0.5+P(0<Z<0.39)"

"=0.5+0.1517"

"=0.65" (approximately)

(ii) We have to find the probability of mean price of 75 selected houses are more than the population mean by RM "1000."

So, "X=(109000+1000)=110000"

"\therefore P(X\geq 110000)= P(Z\geq \frac{110000-109000}{5700})"

"=P(Z\geq 0.18)"

"=0.5-P(0<Z<0.18)"

"=0.5-0.0714"

"=0.43" (approximately)

(iii) "P(107800<X<110100)=P(\frac{107800-109000}{5700}<Z<\frac{110100-109000}{5700})"

"=P(-0.21<Z<0.19)"

"=P(0<Z<0.21)+P(0<Z<0.19)"

"=" "0.0831+0.0753"

"=0.16" (approximately)